An object is moving in a straight line. It initially travels at a speed of 6 meters per second, and it speeds up at a constant acceleration of 4 meters per second each second. The distance d, in meters, that this object travels is given by the equation d = 2t^2 + 6t, where t is in seconds. According to this
equation, how long will it take the object to travel 108 meters?

Answer by Alan3354(69443)

d = 2t^2 + 6t
---------
d = 108
2t^2 + 6t = 108
t^2 + 3t - 54 = 0
(t+9)*(t-6) = 0
t = 6 seconds (ignore the -9 seconds)